I need help for this as soon as possible, thank you.

Winston Inc. is trying to determine the effect of its inventory turnover ratio and days sales outstanding on its cash conversion cycle. Winston's 2015 sales (all on credit) were $177,000 and its cost of goods sold was 75% of sales. It turned over its inventory 8.32 times during the year. Its receivables balance at the end of the year was $13,124.27 and its payables balance at the end of the year was $7,390.99. Using this information calculate the firm's cash conversion cycle. Round your answer to the nearest whole. Round the days amounts in your intermediate calculations to the nearest whole day. Do not round other intermediate calculations.
__ days
Answer #1

Sales =177000
COGS =75% *Sales =75%*177000 =132750
Inventory turnover ratio =8.32

Days Inventory =365/Inventory Turnover =365/8.32 =43.87
Days Sales Outstanding =365/(Sales/Account Receivable) =365/(177000/13124.27)=27.06
Days Account Payable =365/(COGS/Accounts Payable) =365/(132750/7390.99) =20.32

Cash conversion Cycle = Days Inventory Outstanding + Days Sales Outstanding - Days Payable outstanding
=43.87+27.06-20.32 =50.61 Days or 51 days